# Dialogues with AI Reduce Beliefs in Misinformation but Build No Lasting Discernment Skills

### v1

@@ Abstract.
Given the growing prevalence of fake information, including increasingly realistic AI-generated news, there is an urgent need to train people to better evaluate and detect misinformation. While interactions with AI have been shown to durably reduce people’s beliefs in false information, it is unclear whether these interactions also teach people the skills to discern false information themselves. We conducted a month-long study where 67 participants classified news headline-image pairs as real or fake, discussed their assessments with an AI system, followed by an unassisted evaluation of unseen news items to measure accuracy before, during, and after AI assistance. While AI assistance produced immediate improvements during AI-assisted sessions (+21% average), participants’ unassisted performance on new items declined significantly by week 4 (-15.3%). These results indicate that while AI may help immediately, it ultimately degrades long-term misinformation detection abilities.
@@ 1. Introduction
AI Assistants such as ChatGPT, Claude and Grok are increasingly used to evaluate the credibility of online information, from judging the authenticity of news headlines and viral images to answering whether medical claims (Shekar et al., 2024) or political rumors (Kuznetsova et al., 2025) are true. While recent research suggests such systems can reduce belief in specific false claims (Costello et al., 2024), it remains unclear whether these conversations teach humans to detect misinformation or merely shift beliefs about false information with AI assistance.
@@
While both human and machine-based fact-checking methods (Chuai et al., 2024a; Ecker et al., 2010) have been used to annotate and remove false information from online spheres, they struggle with the speed at which misinformation is produced and spread, or fail to be accurate. Additionally, machine learning-based detection algorithms (Wang et al., 2020, 2023; Sha et al., 2023) are computationally expensive and struggle to keep up with the speed at which misinformation can be generated from generative models. Humans are poor at detecting AI-generated content due to the increasing sophistication of generative models and a lack of knowledge about their outputs, such as inconsistent lighting or unnatural textures (Boutadjine et al., 2025). Adding to this challenge, research shows that even when false information is explicitly corrected with human or machine-based corrections (Chuai et al., 2024a; Ecker et al., 2010; Wang et al., 2020, 2023; Sha et al., 2023), people often retain their initial beliefs—a persistent cognitive bias known as the continual influence effect (Johnson and Seifert, 1994a).
While fact-checking systems have shown mixed results on helping people detect misinformation and correct their beliefs about it (Chuai et al., 2024b; Walter et al., 2020), recent work on AI chatbots has demonstrated that dialogues with an AI chatbot can reduce belief in conspiracy theories by 22% (Costello et al., 2024), suggesting potential for AI-assisted belief correction. Furthermore, AI chatbots for education have shown promise in various domains (Cai et al., 2024; Sabnis and Nagashima, 2024; Ceha et al., 2021; Chhibber and Law, 2019), particularly when acting as collaborative partners rather than authoritative sources. This makes the AI chatbot a potential tool to teach people how to detect and update their beliefs about misinformation. Yet emerging research on Human-AI interaction suggests that AI chatbots can foster overreliance (Danry et al., 2023; Kosmyna et al., 2025), with users deferring to system outputs even when incorrect (Danry et al., 2023), potentially undermining the development of independent reasoning skills (Buçinca et al., 2021; Gerlich, 2025). This raises an important question: Do dialogues with AI chatbots actually build lasting misinformation detection skills, or do they create dependency that undermines humans’ independent judgment?
@@
Longitudinal Study: We conduct the first month-long, three-phase longitudinal study measuring misinformation detection effects both with and without AI assistance (N=67).
Dependency Paradox: We demonstrate a critical limitation where AI provides immediate accuracy gains (+21%) but leads to a significant decline in unassisted performance (-15.3% by week 4), particularly for detecting fake content.
@@ 2.1. Why Misinformation Persists
Decades of psychological research explain why misinformation is so resistant to correction once it is initially believed. Classic work on rumor and memory demonstrated that false claims continue to shape reasoning even after being explicitly retracted (Allport and Postman, 1947; Johnson and Seifert, 1994b), a phenomenon now termed the continued influence effect. Meta-analyses and reviews confirm that even detailed corrections rarely eliminate lingering effects (Lewandowsky et al., 2012; Ecker et al., 2010; Walter and Tukachinsky, 2020; Ecker et al., 2022).
Another well-documented mechanism is the illusory truth effect, where repetition increases processing fluency and thus perceived accuracy (Pennycook et al., 2018; Hassan and Barber, 2021; Fazio et al., 2015). Importantly, even prior knowledge does not reliably inoculate against this effect — people with strong factual knowledge are still vulnerable to repeated falsehoods (Fazio et al., 2015). Recent work shows that reminders and retractions can sometimes backfire, particularly when cognitive load is high or attention is divided (Sanderson et al., 2023). To overcome the continued influence effect, simply correcting false information is insufficient. Research demonstrates that people need coherent alternative explanations that fill the explanatory void left behind to effectively remove the influence of misinformation (Ecker et al., 2010; Johnson and Seifert, 1994b; Rich and Zaragoza, 2016; Ithisuphalap et al., 2020).
Together, these findings highlight why misinformation proves so resistant to correction. Heuristics and intuitive reasoning further amplify this vulnerability, as people often rely on fast, effortless judgments rather than deliberate scrutiny (Tversky and Kahneman, 1974; Thompson et al., 2011; Kahneman, 2011). This persistence underscores the need for interventions that not only correct specific falsehoods but also address the psychological mechanisms that sustain them.
@@ 2.2. Interventions to Reduce Misinformation Beliefs
Extensive research has examined how to mitigate misinformation persistence through corrective information provided before, during, or after exposure. Professional fact-checks and platform warning labels have been found to reduce belief in false stories, but often to a modest effect (Chuai et al., 2024a). Even when corrections are well-designed, people frequently continue to draw inferences from the original misinformation, a phenomenon captured by the continued influence effect (Ecker et al., 2010; Rich and Zaragoza, 2016; Walter and Tukachinsky, 2020). Meta-analyses confirm that while fact-checks do exert a positive influence, their success depends heavily on factors such as timing, repetition, context, and cognitive load during exposure (Ecker et al., 2017; Ithisuphalap et al., 2020; Sanderson et al., 2023).
@@ 2.3. AI Dialogues and Persuasion
Dialogue-based interventions differ from static fact-checks by responding to the specific reasons people give for their beliefs. In the strongest demonstration to date, Costello et al. asked participants to describe why they endorsed a conspiracy claim and then instructed the AI chatbot to reply persuasively, but respectfully, with counterarguments tailored to those reasons (Costello et al., 2024). The result was a large and durable reduction in conspiracy belief. Follow-up analyses clarify the mechanism: the dialogues were effective when the AI chatbot provided factual counterevidence and alternative explanations targeted to the user’s arguments; when the system was prompted to persuade without presenting counterevidence, the effect disappeared (Costello et al., [n. d.]). In short, the intervention worked not because an AI was present, but because users experienced having their own arguments acknowledged and then met with “just the facts” that directly undermined those arguments.
@@ 3.2. Data Curation
We randomly selected 55 news items from the MiRAGeNews dataset (Huang et al., 2024), which comprises 12,500 high-quality image-caption pairs—both real news items and AI-generated pairs produced using state-of-the-art generators—along with their ground truth labels (Real or Fake). Figure 4 shows representative examples from the dataset. These selected news items were processed through the architecture discussed in Figure 3 and prompted to output ”Real” or ”Fake”. The output was then compared with the ground truth labels from the MiRAGeNews dataset. The system misclassified 6 items, which we removed from the dataset. The remaining 49 items formed our final dataset, which was distributed across three study phases. Throughout the three phases, participants evaluated 4 randomly selected news items (2 real and 2 fabricated), followed by confidence ratings on 4 previously unseen items to assess whether they could correctly distinguish between real and fake news in unassisted sessions. In phases 2 and 3, the unassisted evaluation sessions included 6 items: 4 novel unseen items and 2 items previously encountered in the preceding phase. This design allowed us to measure both immediate learning on misinformation discernment and retention over time. The items span diverse topics, including elections, protests, court proceedings, diplomatic events, health matters, and disasters, ensuring broad coverage of typical news categories where misinformation commonly appears.
@@ 3.4. Interaction Flow & Conversation Data
Each session follows a standardized protocol. Participants first indicate whether they have seen the news item before, then provide an initial authenticity rating using a binary choice (REAL/FAKE) and confidence slider (0-100%). The system then initiates a conversation where the AI attempts to persuade users toward the correct answer if they were wrong (refer to Figure 13 to view an example from one such conversation). If they were correct, the AI response reinforces the user’s belief, and the user can still interact with the AI system (refer to Figure 12 to view an example from one such conversation).
@@ 6.2. Persuasive AI substantially improved accuracy when assisting the user
6.3. People’s accuracy with AI did not persist when they were unassisted, and their detection abilities declined
Our results show that after interacting with the AI when it is unavailable, their accuracy on unaided items (”After AI support”) declined steadily over the four-week period. When judging new items without AI assistance, accuracy decreased linearly across weeks (, , , ). A direct comparison confirmed that Week 4 accuracy was significantly lower than Week 0 (, ). These results suggest that repeated exposure to AI support improved participants’ judgments with the AI chatbot but eroded their ability to independently identify misinformation when they no longer had access to it.
6.4. Declines in abilities were driven by failures to detect fake content
@@ 7. AI Interaction Strategies and Fake News Discernment
To understand which conversational strategies shaped fake news detection, we applied LLM-as-a-judge (Gu et al., 2024; Zheng et al., 2023) to classify each of the human–AI dialogues into 21 classifier categories. These classifiers capture specific conversational moves such as asking for image forensics, probing deeper, or prompting confidence calibration (see Appendix E.2 for full descriptions). For example, image forensics questions occur when the AI asks about visual artifacts or manipulations in an image (e.g., shadows or lighting), while confidence calibration refers to prompting the user to rate and reflect on how certain they are in their judgment. By coding conversations in this way, we were able to correlate strategy use with accuracy on fake items under both assisted and unassisted conditions. To avoid correlations with rarely occurring strategies, our analysis included only strategies with occurrence rates between 5% and 95% of all conversations with at least 10 positive classifications, and a sufficient variance (). We then calculated Pearson correlation (r) coefficients between each conversational strategy (binary coded by LLM-as-a-judge) and accuracy outcomes under two conditions: AI-assisted accuracy (during Human-AI interaction) and unassisted accuracy (after AI interaction). For each qualifying strategy, we computed correlations separately for fake news items and calculated the difference between unassisted and AI-assisted correlations (diff = - ) to identify whether strategies promoted learning transfer (positive difference) or AI dependence (negative difference). Standard deviations were calculated from the strategy occurrence distributions to assess variability across conversations. An overview of these correlations can be found in Figure 7.
7.1. Strategies paying attention to image artifacts and prior knowledge were associated with higher assisted accuracy
During AI-assisted sessions, strategies that directed attention to concrete cues were the strongest predictors of success. Image forensics questions (, occurring in 22.9% of conversations, ) and recall of prior knowledge (, 68.8%, ) improved assisted accuracy the most. More modest gains came from logic plausibility checks (, 37.5%, ), source checking prompts (, 18.8%, ), and contextual consistency checks (, 33.3%, ). Other strategies had negligible effects (reflect on process, , 43.8%, ) or negative correlations, including confidence calibration (, 43.8%, ), probing deeper across multiple messages (, 91.7%, ), and guiding questions (, 83.3%, ). In short, when the AI helped users look for manipulation or connect back to what they already knew, participants were better able to identify fake content.
7.2. Guiding and probing questions were associated with independent detection abilities
On unaided trials after using the AI, a different pattern emerged. Guiding questions—where the AI in previous interactions had asked users previously to reflect on sources or verifiable details—flipped from negative in assisted sessions () to the strongest predictor of unassisted detection (; diff=, std=, 83.3%, ). Similarly, probing deeper over multiple messages was associated with reduced accuracy while interacting with the AI () but was associated with higher accuracy after interacting with the AI unassisted (; diff=, std=, 91.7%, ). Forensics and knowledge cues again carried over, with questions around image forensics during AI interaction increased both accuracy with the AI and when unassisted afterwards ( with AI vs. unaided; diff=, std=, ). The AI asking the user to recall prior knowledge that is relevant also increased accuracy with and unassisted after interacting with the AI ( with AI vs. unaided; diff=, std=, ).
7.3. Some strategies helped in the moment but undermined learning
Several strategies proved to be only helpful when the AI chatbot was present. Confidence calibration is the clearest case: it slightly reduced accuracy with AI (, 43.8%, ) and strongly undermined unassisted accuracy after AI (, diff=). Similarly, devil’s advocate roleplay provided no assisted benefit (, 12.5%, ) and worsened unaided performance (). These strategies appear to foster dependency by drawing attention to subjective confidence or rhetorical sparring rather than transferable heuristics.
7.4. Balanced strategies supported both assistance and learning
Finally, some conversational moves provided small but consistent benefits across different news items. Image forensics questions (with AI , unassisted after AI , diff=, std=, ) and recall of prior knowledge (with AI , after AI , diff=, std=, ) boosted accuracy regardless of whether the AI was present. Similarly, logic plausibility checks (with AI , after AI , diff=, ) and ask-for-source checking (with AI , without AI , diff=, ) showed modest positive effects in both conditions. These “balanced strategies” offered steady but modest returns, reinforcing their value as design levers for both immediate support and independent discernment.

### v2

@@ Abstract.
Given the growing prevalence of fake information, including increasingly realistic AI-generated news, there is an urgent need to train people to better evaluate and detect misinformation. While interactions with AI have been shown to durably reduce people’s beliefs in false information, it is unclear whether these interactions also teach people the skills to discern false information themselves. We conducted a month-long study where 67 participants classified news headline-image pairs as real or fake, discussed their assessments with an AI system, followed by an unassisted evaluation of unseen news items to measure accuracy before, during, and after AI assistance. While AI assistance produced immediate improvements during AI-assisted sessions (+21% average), participants’ unassisted performance on new items declined significantly by 15.3% in week 4 compared to week 0. These results indicate that while AI may help immediately, it may ultimately degrade long-term misinformation detection abilities.
@@ 1. Introduction
AI Assistants such as ChatGPT, Claude, and Grok are increasingly used to evaluate the credibility of online information, from judging the authenticity of news headlines and viral images to answering whether medical claims (Shekar et al., 2024) or political rumors (Kuznetsova et al., 2025) are true. While recent research suggests such systems can reduce belief in specific false claims (Costello et al., 2024), it remains unclear whether these conversations teach humans to detect misinformation or merely shift beliefs about false information with AI assistance.
@@
While both human and machine-based fact-checking methods (Chuai et al., 2024b; Ecker et al., 2010) have been used to annotate and remove false information from online spheres, they struggle with the speed at which misinformation is produced and spread, or fail to be accurate. Additionally, machine learning-based detection algorithms (Wang et al., 2020, 2023; Sha et al., 2023) are computationally expensive and struggle to keep up with the speed at which misinformation can be generated from generative models. Humans are poor at detecting AI-generated content due to the increasing sophistication of generative models and a lack of knowledge about their outputs, such as inconsistent lighting or unnatural textures (Boutadjine et al., 2025). Adding to this challenge, research shows that even when false information is explicitly corrected with human or machine-based corrections (Chuai et al., 2024b; Ecker et al., 2010; Wang et al., 2020, 2023; Sha et al., 2023), people often retain their initial beliefs—a persistent cognitive bias known as the continual influence effect (Johnson and Seifert, 1994b).
While fact-checking systems have shown mixed results on helping people detect misinformation and correct their beliefs about it (Chuai et al., 2024a; Walter et al., 2020), recent work on AI chatbots has demonstrated that dialogues with an AI chatbot can reduce belief in conspiracy theories by 22% (Costello et al., 2024), suggesting potential for AI-assisted belief correction. Furthermore, AI chatbots for education have shown promise in various domains (Cai et al., 2024; Sabnis and Nagashima, 2024; Ceha et al., 2021; Chhibber and Law, 2019), particularly when acting as collaborative partners rather than authoritative sources. This makes the AI chatbot a potential tool to teach people how to detect and update their beliefs about misinformation. Yet emerging research on Human-AI interaction suggests that AI chatbots can foster overreliance (Danry et al., 2023; Kosmyna et al., 2025), with users deferring to system outputs even when incorrect (Danry et al., 2023), potentially undermining the development of independent reasoning skills (Buçinca et al., 2021; Gerlich, 2025). This raises an important question: Do dialogues with AI chatbots actually build lasting misinformation detection skills, or do they create dependency that undermines humans’ independent judgment?
@@
Longitudinal Study: We conduct the first month-long, three-phase longitudinal study measuring misinformation detection effects both with and without AI assistance (before and after AI use) (N=67).
Dependency/Learning Effects: We demonstrate a trade-off where AI provides immediate accuracy gains (+21%) but leads to a significant decline in unassisted performance after AI interaction (15.3% by week 4 compared to week 0), particularly for detecting fake content; and no significant difference in learning outcomes (unassisted accuracy before AI interaction)
@@ 2.1. Why Misinformation Persists
Decades of psychological research explain why misinformation is so resistant to correction once it is initially believed. Classic work on rumor and memory demonstrated that false claims continue to shape reasoning even after being explicitly retracted (Allport and Postman, 1947; Johnson and Seifert, 1994a), a phenomenon now termed the continued influence effect. Meta-analyses and reviews confirm that even detailed corrections rarely eliminate lingering effects after the first exposure to false information (Lewandowsky et al., 2012; Ecker et al., 2010; Walter and Tukachinsky, 2020; Ecker et al., 2022).
Another well-documented mechanism is repetition of information which increases processing fluency and thus perceived accuracy (also known as the illusory truth effect) (Pennycook et al., 2018; Hassan and Barber, 2021; Fazio et al., 2015). Importantly, even prior knowledge does not reliably inoculate against this effect — people with strong factual knowledge are still vulnerable to repeated falsehoods (Fazio et al., 2015). Recent work shows that reminders and retractions can sometimes backfire, particularly when cognitive load is high or attention is divided (Sanderson et al., 2023). To overcome the continued influence effect, simply correcting false information is insufficient. Instead, research shows that people need coherent alternative explanations that fill the explanatory void left behind to effectively remove the influence of exposure to misinformation (Ecker et al., 2010; Johnson and Seifert, 1994a; Rich and Zaragoza, 2016; Ithisuphalap et al., 2020).
A last relevant mechanism for belief persistence and persuasion in misinformation is the influence of peripheral authority-cues such as professional-looking images or authoritative-seeming sources or explanations (Petty and Cacioppo, 1986; Danry et al., 2025). Experiments show that even nonprobative photographs can inflate the perceived truth of trivia statements, and that adding images to news headlines increases perceived accuracy for both true and false items by boosting familiarity and subjective plausibility (Newman et al., 2012; Smelter and Calvillo, 2020). Related work on deepfakes and AI-generated visuals finds that realistic, evidential-looking images and videos can substantially increase deception, uncertainty, and trust in misleading political content (Groh et al., 2022; Vaccari and Chadwick, 2020; Guo et al., 2025). In low-attention, high-distraction online environments—where people often lack the motivation or cognitive resources to scrutinize claims—such cues can therefore strongly amplify both susceptibility to and persistence of misinformation.
Together, these findings highlight why misinformation proves so resistant to correction. Heuristics and intuitive reasoning further amplify this vulnerability, as people often rely on fast, effortless judgments rather than deliberate scrutiny, which have been found to overwrite the likelihood of such errors happening. (Tversky and Kahneman, 1974; Thompson et al., 2011; Kahneman, 2011). This persistence underscores the need for interventions that not only correct specific falsehoods but also address the psychological mechanisms that sustain them.
@@ 2.2. Interventions to Reduce Misinformation Beliefs
Extensive research has examined how to mitigate misinformation persistence through corrective information provided before, during, or after exposure. Professional fact-checks and platform warning labels have been found to reduce belief in false stories, but often to a modest effect (Chuai et al., 2024b). Even when corrections are well-designed, people frequently continue to draw inferences from the original misinformation, a phenomenon captured by the continued influence effect (Ecker et al., 2010; Rich and Zaragoza, 2016; Walter and Tukachinsky, 2020). Meta-analyses confirm that while fact-checks do exert a positive influence, their success depends heavily on factors such as timing, repetition, context, and cognitive load during exposure (Ecker et al., 2017; Ithisuphalap et al., 2020; Sanderson et al., 2023).
@@ 2.3. AI Dialogues and Persuasion
Dialogue-based interventions differ from static fact-checks by responding to the specific reasons people give for their beliefs. In the strongest demonstration to date, Costello et al. asked participants to describe why they endorsed a conspiracy claim and then instructed the AI chatbot to reply persuasively, but respectfully, with counterarguments tailored to those reasons (Costello et al., 2024). The result was a large and durable reduction in conspiracy belief. Follow-up analyses clarify the mechanism: the dialogues were effective when the AI chatbot provided factual counterevidence and alternative explanations targeted to the user’s arguments; when the system was prompted to persuade without presenting counterevidence, the effect disappeared (Costello et al., ). In short, the intervention worked not because an AI was present, but because users experienced having their own arguments acknowledged and then met with “just the facts” that directly undermined those arguments.
@@ 3.2. Data Curation
We randomly selected 55 news items from the MiRAGeNews dataset (Huang et al., 2024), which comprises 12,500 high-quality image-caption pairs—both real news items and AI-generated pairs produced using state-of-the-art generators—along with their ground truth labels (Real or Fake). Figure 4 shows representative examples from the dataset. These selected news items were processed through the architecture discussed in Figure 3 and prompted to output ”Real” or ”Fake”. The output was then compared with the ground truth labels from the MiRAGeNews dataset. The system misclassified 6 items, which we removed from the dataset. The remaining 49 items formed our final dataset, which was distributed across three study phases. Throughout the three phases, participants evaluated 4 randomly selected news items (2 real and 2 fabricated), followed by confidence ratings on 4 previously unseen items to assess whether they could correctly distinguish between real and fake news in unassisted sessions. In phases 2 and 3, the unassisted evaluation sessions immediately after AI interactions included 6 items: 4 novel unseen items and 2 items previously encountered in the preceding phase. However, we excluded 2 previously seen items from the preceding phase to measure immediate learning on misinformation discernment. The items span diverse topics, including elections, protests, court proceedings, diplomatic events, health matters, and disasters, ensuring broad coverage of typical news categories where misinformation commonly appears.
@@ 3.4. Interaction Flow & Conversation Data
Each session follows a standardized protocol. Participants first indicate whether they have seen the news item before, then provide an initial authenticity rating using a binary choice (REAL/FAKE) and confidence slider (0-100%). The system then initiates a conversation where the AI attempts to persuade users toward the correct answer if they were wrong (refer to Figure 14 to view an example from one such conversation). If they were correct, the AI response reinforces the user’s belief, and the user can still interact with the AI system (refer to Figure 13 to view an example from one such conversation). Our data shows that users changed their initial assessment in 81.14% of conversations when initially incorrect, and maintained their assessment in 98.68% of conversations when initially correct.
@@ 6.1. Descriptive Overview
6.2. Unassisted accuracy before AI interactions showed no to modest declines over time, indicating limited skill learning
To assess whether participants developed improved discernment skills from their AI interactions, we examined baseline accuracy before interacting with the AI in each weekly session (before any AI support). This “Before AI Support” measure provides a cleaner indicator of learning transfer than post-interaction performance, as it captures participants’ independent abilities before any AI interaction. Before AI, accuracy showed a modest decline over the four-week period. Week 4 baseline accuracy was 6.8 percentage points lower than Week 0, though this difference did not reach statistical significance (95% CI: to pp, ). A linear trend analysis similarly indicated a decline of pp per two-week step (, , , ), while Week 2 showed a small non-significant increase ( pp, ). These results suggest that repeated interactions with the AI system did not substantially improve participants’ independent discernment abilities between sessions.
@@ 6.3. Persuasive AI substantially improved accuracy when assisting the user
6.4. People’s accuracy with AI did not persist when they were unassisted after AI interactions, and their detection abilities declined
Our results show that after interacting with the AI when it is unavailable, their accuracy on unaided items (”After AI support”) declined steadily over the four-week period. When judging new items without AI assistance, accuracy decreased linearly across weeks (, , , ). A direct comparison confirmed that Week 4 accuracy was significantly lower than Week 0 (, ).
Within each session, after participants completed their AI-assisted judgments, their accuracy on new unaided items (“After AI Support”) declined more steeply than baseline. When judging items without AI assistance after having used it earlier in the same session, accuracy decreased linearly across weeks (, , , ). Week 4 accuracy was 15.3 percentage points lower than Week 0 (, ). This sharper within-session decline ( pp per step), compared to the more modest between-session baseline change ( pp per step), suggests that the deterioration in “After AI” performance was driven at least in part by short-term dependency or cognitive fatigue effects that emerged immediately following AI use within a session. The fact that baseline performance at the start of subsequent weeks did not show comparable declines indicates some recovery between sessions, though participants did not develop lasting improvements in their independent abilities.
6.5. Declines in unassisted accuracy after AI interactions were driven by failure to detect fake content
@@ 7. AI Interaction Strategies and Fake News Discernment
To examine which conversational strategies shaped fake news detection, we applied an LLM-as-a-judge approach (Gu et al., 2024; Zheng et al., 2023) to classify each human–AI dialogue into 21 strategy categories. These categories capture concrete conversational moves such as asking guided questions, looking for image artifacts, asking to check sources, or asking about their confidence (see Appendix E.2 for full definitions). For example, look for image artifacts refers to moments when the AI directs attention to potential visual manipulation cues (e.g., shadows or lighting inconsistencies), while ask about their confidence prompts users to reflect on how certain they feel about their judgment.
We focus here on fake news items only. Strategies were included in the analysis if they occurred in 5–95% of conversations, had at least 10 positive classifications, and exhibited sufficient variance (). We computed Pearson correlation coefficients between the presence of each strategy (binary coded) and accuracy under three conditions: (1) accuracy before AI assistance (baseline), (2) accuracy during AI-assisted interaction, and (3) unassisted accuracy after interacting with the AI. This allowed us to distinguish strategies associated with in-the-moment performance from those linked to independent detection ability. All reported -values are from two-tailed Pearson tests. Full results for both real and fake news items are reported in Appendix 20.
7.1. Before AI assistance, guided and reflective questioning indexed stronger baseline discernment
Before any AI interaction, several AI interaction strategies were strongly associated with participants’ baseline ability to detect fake news over the 4 weeks. Asking guided questions showed the strongest positive association with pre-AI accuracy (, 83.3%, , ), followed by reflect on progress (, 43.8%, , ) and ask broad questions (, 81.2%, , ). This suggests that reflection and questioning based guidance may lead to learning to better discern between false and true information.
In contrast, several strategies were negatively associated with baseline accuracy. Compare with known facts (, 52.1%, , ), cross verification (, 83.3%, , ), and gave away answer at the start (, 58.3%, , ) were all linked to lower pre-AI performance over the 4 weeks. This suggests that reliance on an AI system that externally validates facts or immediately answers without including the user may lead to the user not learning to discern themselves; or even become worse.
7.2. During AI assistance, perceptual cues helped while reflective prompts hindered performance
During AI-assisted interaction, strategies that directed attention to concrete cues were most consistently associated with improved fake news detection. Look for image artifacts (, 22.9%, ) and user’s prior experiences (, 68.8%, ) showed the largest positive associations with assisted accuracy, although neither reached statistical significance. In contrast, more reflective or dialogically demanding strategies—including asking guided questions (), ask broad questions (), probe deeper over msgs (), and ask about their confidence ()—were associated with reduced performance while the AI was present. The strongest negative association during AI interaction was observed for devils advocate (, 12.5%, , ), suggesting that adversarial framing can substantially disrupt users’ ability to make accurate judgments in the moment.
7.3. After AI interaction, guided questioning and deep probing supported independent detection
A different pattern emerged once participants made judgments after having the AI assistance present. Asking guided questions became the strongest positive predictor of unassisted fake news detection (, 83.3%, , ), despite being negatively associated with accuracy during AI interaction. Similarly, probe deeper over msgs was associated with improved unassisted accuracy (, 91.7%, ). These strategies appear to function as scaffolding mechanisms that slow performance during interaction but support learning and transfer afterward.
Strategies focused on concrete evidence showed smaller but consistent carryover effects. Looking for image artifacts (), relating to the user’s prior experiences (), being asked to check sources (), and checking logic plausibility () were all positively associated with post-AI accuracy, though none reached conventional significance thresholds.
In contrast, being asked about their confidence strongly undermined independent performance right after the interaction (, 43.8%, , ). Devils advocate interactions () and being asked to compare with known facts () also showed negative associations, suggesting that repeatedly challenging users’ confidence or positioning the AI as a corrective authority may weaken users’ trust in their own judgment right after interaction, potentially creating dependency effects.
7.4. Strategies differ in whether they support assistance, learning, or dependence
Taken together, the results reveal a clear dissociation between strategies that help in the moment and those that support learning transfer. Concrete strategies such as image forensics and recalling prior knowledge provide modest benefits both during and after AI use. In contrast, guiding questions and deep probing appear to function as pedagogical strategies: they reduce performance while the AI is present but are associated with stronger independent fake news detection afterward. Finally, confidence calibration and devil’s advocate roleplay consistently undermine unassisted performance, suggesting that these approaches may risk fostering dependence on the AI rather than strengthening users’ own misinformation discernment skills.